Con hno3 is 69% by wight hno3 and has a density of 1.41gm/ml at 20`c.what volume and wght of conc hno3 acid are needed to prepare 100ml of 6M acid?

Grade:12th pass

1 Answers

Vikas TU
7 years ago
 To prepare 100ml of 6M acid => 
moles would be = > 6 x 100/1000 = > 0.6
And 69% by weight,
Volume = mass/density => 69/1.41 = > 48.936 ml
moles => 0.6 = weight/63
Weight = 378/10 = > 37.8 gms.